# From Open Vocabulary to Open World: Teaching Vision Language Models to Detect Novel Objects

### v1

@@ 4.1 Datasets
We evaluate our method with common open world object detection benchmarks used in previous works [72, 33, 58], and proposal a novel benchmark of OWOD for autonomous driving. Common OWOD benchmarks include the superclass-mixed benchmark (M-OWODB) [19] and the superclass-separated benchmark (S-OWODB) [14]. The M-OWODB benchmark combines COCO [27] and PASCAL VOC [8], while the S-OWODB benchmark is based solely on COCO. Both are divided into four distinct tasks, where the model learns some new classes in each task, while the remaining classes are unknown. Additionally, we propose a challenging OWOD benchmark (nu-OWODB) based on nuScenes [2], which consists of real-world driving scenes. The nuScenes dataset captures diverse urban environments, including crowded city streets with many dynamic objects, challenging weather conditions, and dense traffic scenarios with occlusions and complex interactions between agents. In addition, the dataset contains a significant class imbalance, with some classes like cars being much more frequent than others like ambulances or construction vehicles. The nu-OWODB benchmark is divided into three subtasks. Initially, the model is introduced to different types of vehicles. In subsequent tasks, various pedestrians and other traffic participants are introduced. This benchmark simulates the challenges of OWOD in real-world applications.
@@ 4.4 Qualitative Analysis
4.5 Ablation Study
We also try different prompts to estimate the embedding for generic objectness. Tab. 5 shows the ablation experiment on M-OWODB, which shows the word “object” is an appropriate. Note that the generic prompt can be estimated by designing a pretext task, but this will introduce bias towards known classes.
5 Conclusion
Supplementary Material
6 Source Code
The source code of the proposed method is available at https://github.com/343gltysprk/ovow.

### v3

@@ 4.1 Datasets
We evaluate our method with common open world object detection benchmarks used in previous works [75, 34, 59], and propose a novel benchmark of OWOD for autonomous driving. Common OWOD benchmarks include the superclass-mixed benchmark (M-OWODB) [19] and the superclass-separated benchmark (S-OWODB) [14]. The M-OWODB benchmark combines COCO [27] and PASCAL VOC [8], while the S-OWODB benchmark is based solely on COCO. Both are divided into four distinct tasks, where the model learns some new classes in each task, while the remaining classes are unknown. Additionally, we propose a challenging OWOD benchmark (nu-OWODB) based on nuScenes [2], which consists of real-world driving scenes. The nuScenes dataset captures diverse urban environments, including crowded city streets with many dynamic objects, challenging weather conditions, and dense traffic scenarios with occlusions and complex interactions between agents. In addition, the dataset contains a significant class imbalance, with some classes like cars being much more frequent than others like ambulances or construction vehicles. The nu-OWODB benchmark is divided into three subtasks. Initially, the model is introduced to different types of vehicles. In subsequent tasks, various pedestrians and other traffic participants are introduced. This benchmark simulates the challenges of OWOD in real-world applications. For the open vocabulary evaluation, we adopt the LVIS minival [20] benchmark, which is widely used in previous works [70, 29, 5].
@@ 4.3 Results
Quantitative Results on OVD benchmark
Model $AP$ $AP_{r}$ $AP_{c}$ $AP_{f}$
MDETR [20] 24.2 20.9 24.3 24.2
Grounding DINO-T [29] 27.4 18.1 23.3 32.7
DetCLIP-T [70] 34.4 26.9 33.9 36.3
YOLO-World-XL [5] 35.7 26.4 33.9 39.0
Ours 35.7 26.4 33.9 38.9
Table 4: Zero-shot open vocabulary performance
Our method performs OWOD by optimizing class embeddings and additional MSCAL modules, while keeping the parameters adapted from YOLO-World pre-training frozen. As a result, our model maintains the performance of open-vocabulary object detection in a zero-shot manner. Following [5], we evaluate our approach zero-shot on LVIS minival [20] (1203 classes). Known Class Embeddings are initialized from class names, while Pseudo Unknown Embedding is constructed from class names and the “object” prompt. Tab. 4 shows performance comparable to state-of-the-art OVD methods, confirming that our framework unifies OVD and OWOD tasks.
@@ 4.4 Qualitative Analysis
4.5 Ablation Study
Additional ablation studies can be found in the supplemental material.
5 Conclusion
Supplementary Material
6 Additional Ablation Studies
6.1 Generic Objectness Prompt
@@
We tried different prompts to estimate the embedding for generic objectness. Tab. 6 shows the ablation experiment on M-OWODB, which shows the word “object” is an appropriate. We define “object” as the generic prompt in one experiment, then reuse it in all OWOD benchmarks (M-OWODB, S-OWODB, and nu-OWODB) without hand-crafted selection of the prompt across domains. Though it is possible to estimate the generic prompt by designing a pretext task, but this will introduce bias towards known classes.
6.2 Fine-tuning YOLO-World
Method mAP $\uparrow$ U-Recall $\uparrow$ WI $\downarrow$ A-OSE $\downarrow$
YOLO-World (zero-shot) 21.8 2.1 0.0463 12316
YOLO-World (Fine-tuned) 30.0 4.9 0.0419 20039
Ours 28.1 45.5 0.0185 1724
Table 7: Open world performance on nu-OWODB task 1.
Although OVD models perform zero-shot detection by design, we can still fine-tune the model for better performance, especially on datasets from different domains. However, this introduces additional problems in the open-world scenario. As shown in Tab. 7, when we fine-tune the class embeddings of YOLO-World on nu-OWODB, the closed-set performance (mAP) improves, but the model’s resistance to OOD objects is significantly reduced, resulting in a higher A-OSE. In contrast, our method shows significant gains in both known and unknown performance, highlighting the importance of the proposed Pseudo Unknown Embedding and Multi-Scale Contrastive Anchor Learning.
6.3 Ablation on $\alpha$
As shown in Tab. 8, we try different $\alpha$ value varing from 0.2 to 0.8. The result shows that the choice of $\alpha$ does not make a significant impact on performance.
$\alpha$ mAP $\uparrow$ U-Recall $\uparrow$ WI $\downarrow$ A-OSE $\downarrow$
0.2 28.2 38.2 0.0183 1700
0.4 28.1 45.5 0.0185 1724
0.8 27.6 37.8 0.0166 1455
Table 8: Ablation study of $\alpha$ on nu-OWODB task 1.